# On Training Data Influence of GPT Models

### 11 Apr 2024 v1

@@ Abstract
Qingyi ${}^{*\heartsuit}$  Yekun Chaithanks: Equal contribution and shared co-first authorship.       Work done during QL’s internship at ${}^{\spadesuit}$  Shuohuan ${}^{\spadesuit}$ Yu ${}^{\spadesuit}$   Keze ${}^{\heartsuit}$   Hua ${}^{\spadesuit}$$^{\heartsuit}$un Yat-sen University  ${}^{\spadesuit}$aidu Inc. {liuqy95}@mail2.sysu.edu.cn {chaiyekun,wangshuohuan,sunyu02}@baidu.com
@@ 1 Introduction
Furthermore, the majority of this research Pruthi et al. (2020); Guu et al. (2023); K and Søgaard (2021); Koh and Liang (2017); Yeh et al. (2018) has focused on test loss as the primary metric of interest, neglecting other vital performance indicators. Metrics such as BLEU Papineni et al. (2002) and ROUGE Lin (2004) scores are crucial for a thorough evaluation of a model’s capabilities, particularly in the context of generative language models where downstream task performance is paramount.
Additionally, the challenge of generalizability—extending methodologies to accommodate unseen data—persists as a significant barrier Guu et al. (2023). This is particularly critical for models expected to adapt to the dynamic and evolving trajectory of NLP tasks.
@@ 2 Preliminaries
In this section, we revisit the conceptual framework of training data attribution (TDA) methods, aiming to quantify the impact of individual training instances on the performance of models with respect to test data points. We proceed to formally delineate the task of simulating training dynamics.
2.1 Task Definition: Training Dynamics Estimation
@@ 3.3 Featurized Simulation Approach
where $\lambda$ is the discounting factor dictating the degree of L2-regularization, $\hat{\phi}_{t}(\cdot)$ is the test score prediction at step $t$ using Eq.(4).
3.4 Implementing GPTfluence
@@ 4.1.1 GPTDynamics Data Collection
Datasets and GPT Training Scenarios
In subsequent experiments, we refer to the comprehensive training process that employs the aggregated FLAN datasets along with task-specific instructions as instruction tuning. Conversely, the term fine-tuning is reserved to describe the process of individually optimizing models on separate tasks without the use of instructional prompts. Both instruction tuning and fine-tuning processes are encapsulated within our GPTDynamics dataset. We refer to Appendix §A.1 for detailed information.
GPT Backbone
We employed Pythia Biderman et al. (2023), a model suite recently made available to the public, as our foundational architecture. Within this suite, we selected five distinct models based on their sizes, encompassing 14M, 70M, 160M, 410M, 1B, and 2.8B, to ensure a broad range of computational capacities were represented.
@@ 4.1.2 Experiment Setup for Simulators
Baselines
We select TracIn Pruthi et al. (2020), Grad-Dot Charpiat et al. (2019), and Simfluence Guu et al. (2023) as our baselines. Refer to Appendix §A.2 for detailed information.
Evaluation Metrics
We utilize a comprehensive set of metrics, including the Mean Squared Error (MSE) and Mean Absolute Error (MAE) calculated across all training steps, alongside the Spearman correlation coefficient ($\rho$) at the final step, to thoroughly assess performance.
@@ 4.2 Test Loss Estimation
Instruction Tuning
Table 1 presents a comparison between our approach and traditional TDA methods for instruction tuning. GPTfluence demonstrated a distinct edge over Simfluence and other gradient-based TDA techniques across a set of five natural language understanding (NLU) and natural language generation (NLG) tasks, as evidenced by the MSE and MAE metrics for the entire trajectory, alongside the Spearman correlation coefficients at the final time step across various test samples. Examples are shown in Fig. LABEL:fig:loss_a and LABEL:fig:loss_b. Additionally, we observed that while the effectiveness of all evaluated TDA methods in predicting loss trajectories varied with changes in GPT sizes, GPTfluence maintained optimal performance, independent of the GPT scale.
@@
Fine-tuning
In Table 2, it is evident that our approach consistently outperforms Simfluence when it comes to fine-tuning GPT models. On average, our method reduces the MSE and MAE across all training steps by $42\%$ and $28\%$, respectively, when compared to Simfluence. This implies that our method is more robust and adaptable in simulating training dynamics.
@@ 4.3 Generalizing to Test Metric Estimation
Instruction Tuning
We have expanded the evaluation of our model beyond the mere prediction of test loss, now including vital measures such as ROUGE and BLEU scores. This approach is an improvement over previous gradient-based methods that were limited to first-order computations for test loss prediction alone. Our GPTfluence offers more comprehensive applicability. Hence, we evaluated their effectiveness on generative tasks (WebNLG and WMT-16 DE/EN) to validate their generalization capabilities to test metrics.
As for instruction tuning, our findings, displayed in Table 4, demonstrate a superior performance of our method over Simfluence in predicting both BLEU and ROUGE-L scores and for GPTs of varying sizes. Intuitively, We draw some qualitative examples in the Fig. LABEL:fig:bleu_c and LABEL:fig:rougeL_d. Notably, for BLEU simulation on the WMT-16 DE/EN task, as the size of GPT increases, all steps MSE of Simfluence increases, whereas our method maintains a more stable performance, even exhibiting slight improvements from 0.92 to 0.93 in loss prediction accuracy at the final step. This suggests that our model is better equipped to manage more challenging tasks and larger model sizes, leveraging the pre-trained representations and instance interactions.
Fine-tuning
Our method’s superiority remains evident in the fine-tuning scenario, as depicted in Table 4, underscoring the robustness of our feature-based simulation approach. It’s worth noting that the margin by which GPTfluence outperforms Simfluence in BLEU metric simulation is not as pronounced in fine-tuning contexts as it is in instruction tuning settings. This discrepancy is likely due to the richer and more diverse data available in instruction tuning, which accentuates Simfluence’s relative inefficiency, given its independent parameter learning for each training instance and a distinct simulator for each test instance.
@@ 4.4 Ablation Study
Practical Influence via Checkpoints
Our featured simulator is adept at learning from past training dynamics. However, monitoring the training dynamics at every step can be expensive, especially when dealing with large-sized GPTs. Therefore, we conduct experiments to choose training checkpoints at specific intervals to approximate the reality of the neighboring points with the training state of that particular point. Then, we trained our simulator on the approximate training dynamics to find the balance between the cost of collecting training dynamics and the simulator performance.
@@
Empirical Analysis of Markov Order Dependency
Using the first-order Markov process to predict future states based on the prior step, potentially oversimplifies GPT training dynamics. Therefore, we consider the training dynamics as an $n$-th order Markov process ($n=2,3,5,10$) and experiment on both language understanding (RTE) and generative (WebNLG) tasks.
The result can be seen in Fig. LABEL:fig:order. Overall, when considering more preceding training information, the simulation error initially increases and decreases for both datasets, as indicated by the all-steps MSE metric. It suggests that a high order $n$ might introduce noise, leading to a degraded simulator’s performance. Moreover, the final-step Spearman’s $\rho$ shows a significant increase from 0.746 to 0.785 for RTE with the increase of order $n$, but not the same for WebNLG. We guess considering more past training information could improve the prediction accuracy at the final step for more straightforward tasks.
@@
To further explore the impact of various feature representations, we conducted experiments on two types of pre-trained encoders: BERT  [^1] https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 and Pythia  [^2] https://github.com/EleutherAI/pythia with different sizes. Results are shown in Fig. 3. In general, BERT’s feature representations produce better simulation results than the Pythia encoder. This could be due to its ability to encode context information in both directions. Interestingly, we also found that increasing the parameters of the Pythia encoder does not always lead to better performance of the performance simulator.
@@ 4.5 Analysis
Robustness across Varying Model Sizes
We conducted experiments to validate how our simulator handles the complexity of GPTs of different sizes, ranging from 14M to 2.8B, specifically focusing on instruction tuning scenarios.
Results are presented in Fig. 4. Our loss simulation experiments revealed that despite the inconsistent simulation performance trend with increasing GPT size, our featurized model consistently surpassed Simfluence. These findings demonstrate the superiority of our model in effectively capturing and managing model complexity.
@@
Unseen Data Generalization
Unlike Simfluence, which restricts the parameters only indexed by seen samples of past training runs, our GPTfluence can handle unseen samples via sample parameterization. We conducted experiments on RTE and WebNLG tasks in fine-tuning scenarios to further verify the unseen data generalization. For a future training run, we experiment in three different unseen data scenarios: 1) Examples in the training curriculum are unseen; 2) Test examples are unseen; 3) Both examples in training the curriculum and test examples are unseen.
@@
4.6 Use Case: Mislabelled Data Identification
Following previous studies Yeh et al. (2018); Pruthi et al. (2020), we provide the use case for mislabeled data identification to evaluate the TDA method further. We take Pythia-410M as the classifier model, sample a subset of SST-2 as the experimental dataset, and compare GPTfluence with the random selection and TracIn-CP. We refer to Appendix B.3 for a detailed description.
The results are presented in Fig. 5. When evaluating the fraction of mislabel identification, our method exhibits comparable results to a random selection strategy, with a marginal lower performance than TracIn-CP. In contrast, the modest differential should not overshadow the superiority GPTfluence shows on test accuracy. Indeed, GPTfluence surpasses both TracIn-CP and random selection methods, in particular demonstrating the obvious efficacy in detecting mislabeled data at lower fractions of our model when precision is paramount, especially in scenarios with sparse data review.
@@ 5 Related Work
Gradient-Based Approximation Methods
This strand of research capitalizes on gradient information to infer the influence of training instances on model predictions, providing a quantifiable measure of individual data points’ contributions Koh and Liang (2017); Yeh et al. (2018); K and Søgaard (2021). Influence Functions, a pioneering method in this domain, leverages the mathematical framework of influence functions for estimating the impact of dataset perturbations on model predictions. Complementing this, TracIn Pruthi et al. (2020) employs gradient-based approximations to trace the influence of training data on test predictions. Similarly, Grad-Dot Charpiat et al. (2019) uses gradient dot products to approximate the influence of training examples. A contemporary work Xia et al. (2024) that adapts the TracIn framework for models optimized with Adam. LESS incorporates LoRA Hu et al. (2021) and random projection Park et al. (2023) techniques to enhance data selection processes. These methods primarily rely on gradients to quantify data influence, offering tractable solutions with varying degrees of approximation accuracy.
Simulation-Based Approaches
An alternative research vein adopts model-based simulations to represent training dynamics Ilyas et al. (2022); Guu et al. (2023). Simfluence Guu et al. (2023) pioneers the simulation-based category by learning a linear model that predicts the influence of training examples through multiplicative and additive factors, as detailed in §2. Recent efforts Engstrom et al. (2024) have focused on simulating the overall influence of training examples, aiming at predicting the cumulative influence of training data for refined data selection.
@@ A.5 Implementing GPTfluence
We provide the methodology for collecting GPTDynamics data and outline the training process for the featurized simulator within GPTfluence, presented as pseudo-code in Algorithm 1.
@@ B.2 Unseen Data Generalization
We offer in-depth simulations of loss and performance metrics across scenarios involving unseen training data, unseen test data, and both unseen training and test data. Simulation results for fine-tuning are detailed in Table 12, while those for instruction-tuning can be found in Table 12.
@@ B.3 Use Case: Mislabelled Data Identification
We first corrupted the dataset by flipping the label of 40% training samples, and the initial classification accuracy was 0.53. Then, we checked a fraction of the corrupted dataset and fixed mislabels after sorting the dataset using the influence metric, calculated through the TDA method. Finally, we retrained the classifier on the corrected data and reported its test accuracy.
We experiment with the following methods: Random is one of the baselines, where we skip the influence calculation and apply random shuffle [^4] We perform random shuffle ten times with different random seeds and report the average performance.. TracIn-CP takes the self-influence as the influence metric by computing the gradient dot-product of the sample and itself. Similarly, GPTfluence computes the influence metric by simulating multiplicated factor $\alpha$ of the sample to itself.

### 16 Apr 2024 v2

@@ Abstract
Qingyi ${}^{*\heartsuit}$  Yekun Chaithanks: Equal contribution and shared co-first authorship.       Work done during QL’s internship at ${}^{\spadesuit}$  Shuohuan ${}^{\spadesuit}$ Yu ${}^{\spadesuit}$  Qiwei ${}^{\diamondsuit}$  Keze ${}^{\heartsuit}$   Hua ${}^{\spadesuit}$$^{\heartsuit}$un Yat-sen University  ${}^{\spadesuit}$aidu Inc.  ${}^{\diamondsuit}$niversity of Copenhagen {liuqy95}@mail2.sysu.edu.cn {chaiyekun,wangshuohuan,sunyu02}@baidu.com
@@ 1 Introduction
Furthermore, the majority of this research Pruthi et al. (2020); Guu et al. (2023); K and Søgaard (2021); Koh and Liang (2017); Yeh et al. (2018) has focused on test loss as the primary metric of interest, neglecting other vital performance indicators. Metrics such as BLEU Papineni et al. (2002) and ROUGE Lin (2004) scores are crucial for a thorough evaluation of a model’s capabilities, particularly in the context of generative language models where downstream task performance is paramount. Additionally, the challenge of generalizability—extending methodologies to accommodate unseen data—persists as a significant barrier Guu et al. (2023). This is particularly critical for models expected to adapt to the dynamic and evolving trajectory of NLP tasks.
@@ 2 Preliminaries
In this section, we revisit the conceptual framework of training data attribution (TDA) methods, aiming to quantify the impact of individual training instances on the performance of models with respect to test data points.
2.1 Task Definition
@@ 3.2 Featurized Simulation Approach
where $\lambda$ is the discounting factor dictating the degree of L2-regularization, $\hat{\phi}_{t}(\cdot)$ is the test score prediction at step $t$ using Eq.(4). Refer to Algorithm 1 for the pseudo-code.
@@ 4.1.1 GPTDynamics Data Collection
Datasets and GPT Training Scenarios  In subsequent experiments, we refer to the comprehensive training process that employs the aggregated FLAN datasets along with task-specific instructions as instruction tuning. Conversely, the term fine-tuning is reserved to describe the process of individually optimizing models on separate tasks without the use of instructional prompts. Both instruction tuning and fine-tuning processes are encapsulated within our GPTDynamics dataset. We refer to Appendix §A.1 for detailed information.
GPT Backbone  We employed Pythia Biderman et al. (2023), a model suite recently made available to the public, as our foundational architecture. Within this suite, we selected five distinct models based on their sizes, encompassing 14M, 70M, 160M, 410M, 1B, and 2.8B, to ensure a broad range of computational capacities were represented.
@@ 4.1.2 Experiment Setup for Simulators
Baselines  We select TracIn Pruthi et al. (2020), Grad-Dot Charpiat et al. (2019), and Simfluence Guu et al. (2023) as our baselines. Refer to Appendix §A.2 for detailed information.
Evaluation Metrics  We utilize a comprehensive set of metrics, including the Mean Squared Error (MSE) and Mean Absolute Error (MAE) calculated across all training steps, alongside the Spearman correlation coefficient ($\rho$) at the final step, to thoroughly assess performance.
@@ 4.2 Test Loss Estimation
Instruction Tuning  Table 1 presents a comparison between our approach and traditional TDA methods for instruction tuning. GPTfluence demonstrated a distinct edge over Simfluence and other gradient-based TDA techniques across a set of five natural language understanding (NLU) and natural language generation (NLG) tasks, as evidenced by the MSE and MAE metrics for the entire trajectory, alongside the Spearman correlation coefficients at the final time step across various test samples. Examples are shown in Fig. LABEL:fig:loss_a and LABEL:fig:loss_b. Additionally, we observed that while the effectiveness of all evaluated TDA methods in predicting loss trajectories varied with changes in GPT sizes, GPTfluence maintained optimal performance, independent of the GPT scale.
@@
Fine-tuning  In Table 2, it is evident that our approach consistently outperforms Simfluence when it comes to fine-tuning GPT models. On average, our method reduces the MSE and MAE across all training steps by $42\%$ and $28\%$, respectively, when compared to Simfluence. This implies that our method is more robust and adaptable in simulating training dynamics.
@@ 4.3 Generalizing to Test Metric Estimation
We have expanded the evaluation of our model beyond the mere prediction of test loss, now including vital measures such as ROUGE and BLEU scores. We have not reported the performance of TracIn and Grad-Dot baselines due to its inability on such metric predictions.
Instruction Tuning  As for instruction tuning, our findings, displayed in Table 4, demonstrate a superior performance of our method over Simfluence in predicting both BLEU and ROUGE-L scores and for GPTs of varying sizes. Intuitively, We draw some qualitative examples in the Fig. LABEL:fig:bleu_c and LABEL:fig:rougeL_d. Notably, for BLEU simulation on the WMT-16 DE/EN task, as the size of GPT increases, all steps MSE of Simfluence increases, whereas our method maintains a more stable performance, even exhibiting slight improvements from 0.92 to 0.93 in loss prediction accuracy at the final step. This suggests that our model is better equipped to manage more challenging tasks and larger model sizes, leveraging the pre-trained representations and instance interactions.
Fine-tuning Our method’s superiority remains evident in the fine-tuning scenario, as depicted in Table 4, underscoring the robustness of our feature-based simulation approach. It’s worth noting that the margin by which GPTfluence outperforms Simfluence in BLEU metric simulation is not as pronounced in fine-tuning contexts as it is in instruction tuning settings. This discrepancy is likely due to the richer and more diverse data available in instruction tuning, which accentuates Simfluence’s relative inefficiency, given its independent parameter learning for each training instance and a distinct simulator for each test instance.
@@ 4.4 Ablation Study
Practical Influence via Checkpoints  Our featured simulator is adept at learning from past training dynamics. However, monitoring the training dynamics at every step can be expensive, especially when dealing with large-sized GPTs. Therefore, we conduct experiments to choose training checkpoints at specific intervals to approximate the reality of the neighboring points with the training state of that particular point. Then, we trained our simulator on the approximate training dynamics to find the balance between the cost of collecting training dynamics and the simulator performance.
@@
Empirical Analysis of Markov Order Dependency  Using the first-order Markov process to predict future states based on the prior step, potentially oversimplifies GPT training dynamics. Therefore, we consider the training dynamics as an $n$-th order Markov process ($n=2,3,5,10$) and experiment on both language understanding (RTE) and generative (WebNLG) tasks.
The result can be seen in Fig. LABEL:fig:order. Overall, when considering more preceding training information, the simulation error initially increases and decreases for both datasets, as indicated by the all-steps MSE metric. It suggests that a high order $n$ might introduce noise, leading to a degraded simulator’s performance. Moreover, the final-step Spearman’s $\rho$ shows a significant increase from 0.746 to 0.785 for RTE with the increase of order $n$, but not the same for WebNLG. We guess considering more past training information could improve the prediction accuracy for NLU tasks.
@@
Impact of Different Feature Representations To further explore the impact of various feature representations, we conducted experiments on two types of pre-trained encoders: BERT  [^1] https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 and Pythia  [^2] https://github.com/EleutherAI/pythia with different sizes. Results are shown in Fig. 3. In general, BERT’s feature representations produce better simulation results than the Pythia encoder. This could be due to its ability to encode context information in both directions. Interestingly, we also found that increasing the parameters of the Pythia encoder does not always lead to better performance of the performance simulator.
@@ 4.5 Analysis
Robustness across Varying Model Sizes  We conducted experiments to validate how our simulator handles the complexity of GPTs of different sizes, ranging from 14M to 2.8B, specifically focusing on instruction tuning scenarios. Results are presented in Fig. 4. Our loss simulation experiments revealed that despite the inconsistent simulation performance trend with increasing GPT size, our featurized model consistently surpassed Simfluence. These findings demonstrate the superiority of our model in effectively capturing and managing model complexity.
@@
Unseen Data Generalization  Unlike Simfluence, which restricts the parameters only indexed by seen samples of past training runs, our GPTfluence can handle unseen samples via sample parameterization. We conducted experiments on RTE and WebNLG tasks in fine-tuning scenarios to further verify the unseen data generalization. For a future training run, we experiment in three different unseen data scenarios: 1) Examples in the training curriculum are unseen; 2) Test examples are unseen; 3) Both examples in training the curriculum and test examples are unseen.
@@ 5 Related Work
Gradient-Based Approximation Methods  This strand of research capitalizes on gradient information to infer the influence of training instances on model predictions, providing a quantifiable measure of individual data points’ contributions Koh and Liang (2017); Yeh et al. (2018); K and Søgaard (2021). Influence Functions, a pioneering method in this domain, leverages the mathematical framework of influence functions for estimating the impact of dataset perturbations on model predictions. Complementing this, TracIn Pruthi et al. (2020) employs gradient-based approximations to trace the influence of training data on test predictions. Similarly, Grad-Dot Charpiat et al. (2019) uses gradient dot products to approximate the influence of training examples. A contemporary work Xia et al. (2024) that adapts the TracIn framework for models optimized with Adam. LESS incorporates LoRA Hu et al. (2021) and random projection Park et al. (2023) techniques to enhance data selection processes. These methods primarily rely on gradients to quantify data influence, offering tractable solutions with varying degrees of approximation accuracy.
Simulation-Based Approaches  An alternative research vein adopts model-based simulations to represent training dynamics Ilyas et al. (2022); Guu et al. (2023). Simfluence Guu et al. (2023) pioneers the simulation-based category by learning a linear model that predicts the influence of training examples through multiplicative and additive factors, as detailed in §2. Recent efforts Engstrom et al. (2024) have focused on simulating the overall influence of training examples, aiming at predicting the cumulative influence of training data for refined data selection.
@@ B.2 Unseen Data Generalization
We offer in-depth simulations of loss and performance metrics across scenarios involving unseen training data, unseen test data, and both unseen training and test data. Simulation results for fine-tuning are detailed in Table 12, while those for instruction-tuning can be found in Table 12. Illustration examples are shown in the Fig. LABEL:fig:unsen_ft.
@@ B.3 Use Case: Mislabelled Data Identification
Following previous studies Yeh et al. (2018); Pruthi et al. (2020), we provide the use case for mislabeled data identification to evaluate the TDA method further.
Experimental Setting
We take Pythia-410M as the classifier model and sample a subset of SST-2 as the experimental dataset. We experiment with the following methods: Random is one of the baselines, where we skip the influence calculation and apply random shuffle [^4] We perform random shuffle ten times with different random seeds and report the average performance.. TracIn-CP takes the self-influence as the influence metric by computing the gradient dot-product of the sample and itself. Similarly, GPTfluence computes the influence metric by simulating multiplicated factor $\alpha$ of the sample to itself. The results are shown in the Fig. 5
Results
The results are presented in Fig. 5. When evaluating the fraction of mislabel identification, our method exhibits comparable results to a random selection strategy, with a marginal lower performance than TracIn-CP. In contrast, the modest differential should not overshadow the superiority GPTfluence shows on test accuracy. Indeed, GPTfluence surpasses both TracIn-CP and random selection methods, in particular demonstrating the obvious efficacy in detecting mislabeled data at lower fractions of our model when precision is paramount, especially in scenarios with sparse data review. We first corrupted the dataset by flipping the label of 40% training samples, and the initial classification accuracy was 0.53. Then, we checked a fraction of the corrupted dataset and fixed mislabels after sorting the dataset using the influence metric, calculated through the TDA method. Finally, we retrained the classifier on the corrected data and reported its test accuracy.